OpenStudy (anonymous):

In a survey of 400 teenage shoppers at a large mall, 240 said they shopped at Abernathy's, 210 said they shopped at Bongo Republic, and 90 said they shopped at both stores. How many of the teenage shoppers surveyed did not shop at either store?

OpenStudy (anonymous):

400 shoppers 240 shop at A = ONLY at A + both A & B 210 shop at B = ONLY at B + both A & B total amount of people so far = 240+210 - both A&B = 450-90 = 360 Amount of people left = 40